An investment center manager is evaluating two projects (Project 1 and Project 2). The firm's required rate of return is 15%. The manager's current ROI is 20%. These two projects are not mutually exclusive. (Project data is expressed in present value.)

Project 1 will provide $185,000 in earnings and require investment of $800,000.

Project 2 will provide $105,000 in earnings and require investment of $565,000.

Which of the following is true?

Answers: a.

Using ROI, the manager is likely to reject Project 1 and accept Project 2, even though both have positive residual income.

b.

Using either ROI or residual income, the manager is likely to accept Project 1 and reject Project 2.

c.

Using either ROI or residual income, the manager is likely to reject Project 1 and accept Project 2.

d.

Using ROI, the manager is likely to accept Project 1 and reject Project 2, even though both have positive residual income.
Answer #1
Project 1:
ROI 23% =185000/800000
Residual income 65000 =185000-(800000*15%)
Project 2:
ROI 19% =105000/565000
Residual income 20250 =105000-(565000*15%)
ROI or residual income both are higher for Project 1
Using either ROI or residual income, the manager is likely to accept Project 1 and reject Project 2.
Option B is correct